OpenStudy (anonymous):

Use formulas to find the surface area of a right prism with a rectangular base.The sides of the base of the prism are 2 m by 6 m, and the high of the prism is 5 m.

OpenStudy (jhannybean):

Did you see how i found the surface area in the last problem you posted?

OpenStudy (jhannybean):

You can use that method to find this surface area.

OpenStudy (anonymous):

I dont understand. what do I plug in ? @Jhannybean

OpenStudy (jhannybean):

Surface area of Right prism  \[\large 2(w\cdot l) + 2(w\cdot h) + 2(l\cdot h) = 2[(w\cdot l) + (w\cdot h) + (l\cdot h)]\]\[\large l = 6 \ m \ \ , \ w = 2 \ m \ , \ h=5 \ m \] plug all these values in for the simplified equation and you will find your surface area :)